# On the foundation of quantum decision theory

### v2

@@ Abstract
Quantum decision theory is introduced here, and new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, next on arguments for the Born rule, that is the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that each accessible variable can be seen as a function of a specific inaccessible variable. Another assumption is that there exist two maximal accessible decision processes in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example. Finally, a broad discussion of decision processes is given.
Keywords: Accessible decision variabel; action; Born rule; decision; Hilbert space formalism; illustrating example; quantum probabilities.
@@ 1 Introduction
Political decisions are important, and may have large consequences. In an autocratic country political decisions are made by sovereign leaders, having gained their positions through power exchanges. In democratic countries, at least in principle people may control their leaders’ decisions.
In this article I will concentrate on decisions made by single persons. From a psychological point of view, such decisions have been thoroughly reviewed by Newell et al. (2015). The book emphasizes the relationship between learning and decisions, arguing that the best way to understand how and why decisions are made is in the context of learning and knowledge acquisition which precedes them, and the feedback which follows. Some of the discussions in op. cit. are also relevant for the developments in the present article.
@@
A person always has his history, and through his history he learns how to handle decisions. Hence learning is important. A person’s learning history starts when he or she is a child, when his or her ideals usually are the parents. Later, he or she is influenced by other people, but we can never disregard such influences as a background for a person’s decisions, even in cases where the person claims that his aim is just to optimize his behaviour according to some criterion.
In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory. The argument will be further elucidated later in this article.
Admittedly, the decision foundation based upon optimizing expected utility has had and has still a strong theoretical support. There are several axiom systems that can motivate such a basis, and the mathematics behind such axioms can be traced back to Bernoulli and Pascal. Newell et al. (2015) discuss one such axiom, the sure thing principle: If a person prefers option A to option B under the condition X, and he also prefers A to B under the opposite condition not-X, then he always prefers A to B.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory (QDT). This theory is completely based upon quantum mechanics, whose foundation is discussed in the book Helland (2021) and in some newer articles, Helland (2022a,b,c, 2023a,b,c, 2024a).
Pionering articles on quantum decision theory are Aerts (2009) and Yukalov and Sornette (2014); see also the book by Busemeyer and Bruza (2012). More recent review articles are Ashtiani and Azgomi (2015) and Pothos an Busemeyer (2022). My approach will be independent of this literature; a completely new entrance to the theory will be aimed at.
The plan of this article is as follows: I first review briefly a new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications is then given, and used to illustrate some of the peculiarities of quantum probabilities. Then my foundation of QDT is further discussed from several points of view before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b,c 2023a,b,c,2024a). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,...$ are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
@@
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take ϕ=italic-$\phi=$ (position, momentum) and ϕ=italic-$\phi=$ full spin vector. In the last example, say in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$.
Here these variables are theoretical variables coupled to some physical situation. If necessary, an observed variable can be modeled as a theoretical variable plus some random error. Following Zwirn (2016, 2020), every description of reality must be seen from the point of view of some observer. Hence we can assume that the variables also exist relative to $C$.
@@
But observers may communicate. The mathematical model developed in the articles above is equally valid relative to a group of people that can communicate about the physics and about the various theoretical variables. This gives a new version of the theory, a version where all theoretical variables are defined jointly for such a group. The only difference here is that, for the variables to function during the communication, they must always be possible to define them in words.
@@
In the above two examples there are also maximal accessible variables: In the first example either position or momentum, in the second example the spin component $\theta^{a}$ in some direction $a$. From a mathematical point of view, an accessible variable $\theta$ is called maximal if there is no other accessible variable $\lambda$ such that $\theta=f(\lambda)$ for some non-invertible function $f$. In general, the term ‘maximal’ will then be seen to be maximal with respect to the partial ordering of variables given by $\alpha\leq\beta$ iff $\alpha=f(\beta)$ for some function $f$.
@@
Two different acccessible variables $\theta$ and $\eta$ are said to be related if there is a transformation $k$ in $\phi$-space and a function $f$ such that $\theta=f(\phi)$ and $\eta=f(k\phi)$. In Helland (2024a) it is shown that if both $\theta$ and $\eta$ are maximal and take $r$ values, then $\phi$ can be constructed in a natural way such that there is an accessible variable $\lambda$ that is in one-to-one correspondence with $\theta$, and such that $\lambda$ and $\eta$ are related according to this definition.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ even when they are different maximal accessible variables. Variables that are maximal, different and not in one-to-one correspondence, are called - following Niels Bohr - for complementary.
@@
In the present paper, we will concentrate on accessible variables that take a finite number of values, say $r$. Then it can be shown that pairs of maximal accessible variables are either related, are in one-to-one correspondence or both. The following Theorem is proved in Helland (2024a):
Theorem 1 Assume that there exist two different maximal accessible variables $\theta$ and $\eta$, each taking $r$ values, and not in one-to-one correpondence. Then, there exists an $r$-dimensional Hilbert space $\mathcal{H}$ describing the situation, and every accessible variable in this situation will have an associated self-adjoint operator in $\mathcal{H}$.
@@
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues of the corresponding operator are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$.
If $\theta$ is maximal as an accessible variable, the eigenvectors of $A^{\theta}$ have a similar interpretation. In my opinion, such eigenvectors, where $\theta$ is some meaningful variable, should be tried to be taken as the only possible state vectors. These have straightforward interpretations, and from this version of the theory, also a number of so-called ‘quantum paradoxes’ can be illuminated, see Helland (2023a,c).
This not only points at a new foundation of quantum theory, and it also suggests a general epistemic interpretation of the theory: Quantum theory is not directly a theory about the world, but a theory about an actor’s knowledge of the world. Versions of such an interpretation already exist, and they are among the very many suggested interpretations of quantum mechanics.
@@
To any ket vektor $|a;k\rangle$ there corresponds a bra vector $\langle a;k|$, the corresponding complex conjugate row vector. Scalar products between two state vectors can then be written $\langle a;k|b;j\rangle$. I will also need scalar products of the form $\langle a;k|A|a;k\rangle$, where $A$ is an operator.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he or she has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values $1,...,r$, such that $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$. If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in relation to a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to take a decision. Let two different maximal decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$, and $\eta=j$ corresponds to the action $b_{j}$ $(j=1,...,r)$. Then, by the theory of the previous subsection, we can model the situation by using quantum theory.
Note that both $\theta$ and $\eta$ may be vector variables. Say that $\theta=(\theta_{1},...,\theta_{n})$, where each $\theta_{j}$ is a simpler decision variable. Then this corresponds to a situation where $C$, in addition to the difficult decision given by $\eta$, is faced with $n$ more simple decisions. Such a situation is not uncommon. In each situation where we shall make a difficult decision, we will be in a context where also a number of trivial decisions may have to be made just in order to survive and to function well in the given context. For many people, these trivial decision occupy a large portion of their mind, such a large part that the vector decision variable $\theta$ also must be considered to be maximal. The assumption that both $\theta$ and $\eta$ take the same number of values $r$, can be satisfied by artificially adding some actions to one of the decision problems.
The simple model above does not cover all situations. Sometimes we have a choice between an infinite number of possibilities, and sometimes the outer context changes during the decision process. Nevertheless, the simple model is a good starting point.
In this article, I will mainly focus on maximal decision problems. A thorough discussion including decisions that are not maximal, is given in Mogiliansky et al. (2009).
@@ 2.3 The Hilbert space formalism for decisions
It is well known that our minds may be limited, for instance when faced with difficult decisions. I will first mention a side result in this direction from the present development.
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneous have in mind any other maximal accessible varable which is related to $\theta$, but not related to $\eta$. It was claimed in Helland (2022c) that the violation of a famous inequality by practical Bell experiments, can be understood on the basis of this theorem. See also Helland (2023b), where a corresponding theorem is formulated without any symmetry assumption.
Note that this result has the qualification ‘at the same time’, and indicate a specific restriction to two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
For the present article, however, the direct results from Helland (2022b, 2024a) are equally important. Consider again a decision situation, and assume the simple model of the previous Subsection. In particular let $C$ at the same time be confronted with at least two different maximal related decision processes. Then the following hold:
- Each decision variable $\eta$ is associated with a self-adjoint operator $A^{\eta}$, whose eigenvalues are the possible values of $\eta$.
@@
- In the maximal case, the eigenvectors of the operator can be given interpretations: They are coupled to one particular decision process and a specific choice in this decision process: In concrete terms, the eigenvectors $v$ are in one-to-one correspondence with 1) some maximal accessible decision variable $\eta$, and 2) a specific value $u$ of $\eta$. In other words, the possible eigenvectors are in one-to-one correspondence with 1) the question ‘Which decision process?’ and 2) ‘Which action did this decision process lead to?’.
@@
This can be taken as a starting point of quantum decision theory, but to develop this theory further, we need to be able to calculate probabilities for the various decisions.
@@ 2.4 The Born rule; general arguments
where the pure state $|b;j\rangle$ corresponds to the event $\theta^{b}=u^{b}_{j}$, and where $\mathbf{u}^{b}=(u^{b}_{1},...,u^{b}_{r})$ is a vector of actual or potential values of the parameter $\theta^{b}$.
The interpretation of the likelihood effect $F^{b}(\mathbf{u}^{b};z^{b},\tau)$ can be formulated as follows: (1) We have posed some inference question on the accessible conceptual variable $\theta^{b}$. (2) We have specified the relevant likelihood for the data. The question itself and the likelihood for all possible answers of the question, perhaps formulated in terms of state vectors, can be recovered from the likelihood effect.
The likelihood effect is closely connected to the concept of an operator-valued measure; see a discussion in Helland (2021). Since the focused question here assumed discrete data, each likelihood is in the range $0\leq p\leq 1$. In the quantum mechanical literature, an effect is any operator with eigenvalues in the range $[0,1]$.
@@
Assumption 1 Consider in the context $\tau$ an epistemic setting where the the likelihood principle from statistics is satisfied, and the whole situation is acted upon by an experimentalist $C$ whose decisions can be modelled to be influenced by a superior being $D$. Assume that $D$’s relevant probabilities for the situation are given by the numbers $q$, and that $D$ can be seen to be perfectly rational in agreement with the Dutch Book Principle.
@@
A situation where Assumption 1 holds will be called a rational epistemic setting. It will be seen to imply essential aspects of quantum mechanics. Below we will discuss whether or not it also can be coupled to certain macroscopic situations, in particular decision situations.
In Helland (2021), a generalized likelihood principle is proved from the ordinary likelihood principle: Given some experiement, or more generally, some context $\tau$ connected to an experiment, any experimental evidence will under weak assumptions be a function of the likelihood effect $F$. In particular, the relevant probabilities $q$ sre functions of $F$: $q(F|\tau)$.
Theorem 2 Assume a rational epistemic setting, and assume a fixed context $\tau$. Let $F_{1}$ and $F_{2}$ be two likelihood effects in this setting, and assume that $F_{1}+F_{2}$ also is an effect. Then the experimental evidences, taken as the epistemic probabilities related to the data of the performed experiments, satisfy
@@
The further derivations rely on a very elegant recent theorem by Busch (2003): Let in general $\mathcal{H}$ be any separable Hilbert space. Recall that an effect $F$ is any operator on the Hilbert space with eigenvalues in the range $[0,1]$. A generalized probability measure $\mu$ is a function on the effects with the properties
@@
It is now easy to see that $q(F|\tau)$ on the likelihood effects of the previous Section is a generalized probability measure if Assumption 1 holds: (1) follows since $q$ is a probability; (2) since $F=I$ implies that the likelihood is 1 for all values of the theoretical variable; finally (3) is a consequence of the corollary of Theorem 2. Hence there is a density operator $\rho=\rho(\tau)$ such that $q(F|\tau)=\mathrm{trace}(\rho(\tau)F)$ for all likelihood effects $F=F(\mathbf{u};z,\tau)$. This is a result that is valid for all experiments, and it can be seen as a first general version of Born’s formula.
@@
Define now a perfect experiment as one where the measurement uncertainty can be disregarded. The quantum mechanical literature operates very much with perfect experiments which result in well-defined states $|j\rangle$. From the point of view of statistics, if, say the 99% confidence or credibility region of $\theta^{b}$ is the single point $u_{j}^{b}$, we can infer approximately that a perfect experiment has given the result $\theta^{b}=u_{j}^{b}$.
@@
An advantage of using the version of Gleason’s theorem due to Busch (Theorem 3) in the derivation of Born’s formula, is that this version is valid also in dimension 2. Other derivations using the same point of departure, are Caves et al. (2004) and Wright and Weigert (2019). In Wright and Weigert (2021) the class of general probabilistic theories which also admit Gleason-type theorems is identified. But for instance, Auffeves and Granger (2019) derive the Born formula from other postulates.

### v3

@@ Abstract
Quantum decision theory is introduced here, and a new basis for this theory is proposed. It is first based upon the author’s general arguments for the Hilbert space formalism in quantum theory, and next on arguments for the Born rule, which is the basis for calculating quantum probabilities. A basic notion behind the quantum theory foundation is that of theoretical variables, that are divided into accessible and inaccessible ones. This is here specialized to decision variables. It is assumed that each accessible variable can be seen as a function of a specific inaccessible variable. Another assumption is that there exist two maximal accessible decision processes in the given situation. Two basic assumptions behind the Born rule are 1) the likelihood principle, 2) the actor in question has motivations that can be modeled by a hypothetical perfectly rational higher being. The theory is illustrated by a medical example.
Keywords: Accessible decision variable; action; Born rule; decision; Hilbert space formalism; illustrating example; quantum probabilities.
@@ 1 Introduction
Political decisions are important, and may have large consequences. In an autocratic country political decisions are made by sovereign leaders, who have gained their positions through power exchanges. In democratic countries, at least in principle people may control their leaders’ decisions.
In this article we will concentrate on decisions made by single persons. From a psychological point of view, such decisions have been thoroughly reviewed by Newell et al. (2015). The book emphasizes the relationship between learning and decisions, arguing that the best way to understand how and why decisions are made is in the context of learning and knowledge acquisition that precedes them, and the feedback that follows. Some of the discussions in op. cit. are also relevant for the developments in the present article.
@@
A person always has his history, and through his history he learns how to handle decisions. Hence learning is important. In standard economic theory and statistical theory, the problem of optimizing behaviour is reduced to optimizing expected reward or utility. I will claim that this does not quite correspond to the way people act during a decision process. This statement can be supported by a number of empirical investigations; see for instance the review articles mentioned below in connection to quantum decision theory.
Admittedly, the decision foundation based upon optimizing expected utility has had and has still strong theoretical support. There are several axiom systems that can motivate such a basis, and the mathematics behind such axioms can be traced back to Bernoulli and Pascal. Newell et al. (2015) discuss one such axiom, the sure thing principle: If a person prefers option A to option B under the condition X, and he also prefers A to B under the opposite condition not-X, then he always prefers A to B.
@@
Theories of decisions can take many points of departure. In this article we will describe a new foundation of a comparatively recent theory: Quantum decision theory. This theory is completely based upon quantum mechanics, whose foundation is discussed in the book Helland (2021) and in some newer articles Helland (2022a,b, 2023a,b).
Pionering articles on quantum decision theory are Aerts (2009) and Yukalov and Sornette (2014); see also the book by Busemeyer and Bruza (2012). More recent review articles are Ashtiani and Azgomi (2015) and Pothos an Busemeyer (2022). Our approach will be independent of this literature; a completely new entrance to the theory will be aimed at.
The plan of this article is as follows: I first review briefly a new approach to the foundation of quantum theory, and then I use this to derive and motivate a new foundation of quantum decision theory, taking as a basis a simple model of a person’s decisions. First, the Hilbert space formalism of decision variables is derived. Then the Born formula, the basis for computing quantum probabilities, is argued for under weak assumptions, and this is specialized to a decision setting. An example of applications are then given, and used to illustrate some of the peculiarities of quantum probabilities. Then the interpretation is briefly discussed before I give some concluding remarks.
@@ 2.1 On the foundation of quantum theory in general
A completely new approach towards quantum foundations is proposed in Helland (2021, 2022a,b, 2023a,b,c,2024a). The basis of this approach is an observer who is in some physical situation. In this situation there are theoretical variables, and certain of these variables, say $\theta,\lambda,\eta,...$ are related to the observer $C$. Some of the variables are accessible to him, meaning roughly that it is, in some future, in principle possible to obtain as accurate values as he wishes on the relevant variable. Others are inaccessible. Examples of the latter are the vector (position, momentum) of a particle at some time, or the full spin vector of a spin particle, an imagined vector whose discretized projection in the direction $a$ is the spin component in that direction. The terms ‘accessible’ and ‘inaccessible’ are primitive notions of the theory.
@@
The main assumption of my theory is then as follows: Related to an observer $C$ there is an inaccessible variable $\phi$ such that all accessible variables can be seen as functions of $\phi$. In the two examples above one can take ϕ=italic-$\phi=$ (position, momentum) and ϕ=italic-$\phi=$ full spin vector. In the last example, in the spin 1/2 case, one can model the discrete spin component $\theta^{a}$ in direction $a$ as $\mathrm{sign}(\mathrm{cos}(a,\phi))$. Giving $\phi$ a reasonable distribution here, results in a correct distribution of each $\theta^{a}$. Here these variables are theoretical, variables coupled to some physical situation. Following Zwirn (2016, 2020), every description of reality must be seen from the point of view of some observer. Hence we can assume that the variables also exist relative to $C$.
@@
But observers may communicate. The mathematical model developed in the articles above is equally valid relative to a group of people who can communicate about physics and the various theoretical variables. This gives a new version of the theory, a version where all theoretical variables are defined jointly for such a group. The only difference here is that, for the variables to function during the communication, it must always be possible to define them in words.
@@
In the above two examples there are also maximal accessible variables: In the first example either position or momentum, in the second example the spin component $\theta^{a}$ in some direction $a$. In general, an accessible variable $\theta$ is maximal if there is no other accessible variable $\lambda$ such that $\theta=f(\lambda)$ for some non-invertible function $f$.
@@
Two different accessible variables $\theta$ and $\eta$ are said to be related if there is a transformation $k$ in $\phi$-space and a function $f$ such that $\theta=f(\phi)$ and $\eta=f(k\phi)$. In Helland (2023b) it is shown that if both $\theta$ and $\eta$ are maximal and take $r$ values, then $\phi$ can be constructed in a natural way, there is an accessible variable $\lambda$ that is in one-to-one correspondence with $\theta$, all this such that $\lambda$ and $\eta$ are related according to this definition.
@@
In this sense, one can always think of some relation between $\theta$ and $\eta$ when they are different maximal accessible variables. Such variables are called - following Niels Bohr - for complementary.
@@
In the present paper, we will concentrate on accessible variables that take a finite number of values, say $r$. Then the following first theorem is proved in the above papers: Given a situation with two different maximal accessible variables $\theta$ and $\eta$, there corresponds to every accessible variable a self-adjoint operator in some complex Hilbert space $\mathcal{H}$. With ‘different’ I here mean really different, not two maximal variables that are one-to-one functions of each other. And a ‘self-adjoint operator’ is just a complex $r\times r$ matrix with real eigenvalues. In particular then, there is such an operator $A^{\theta}$ corresponding to $\theta$ and an operator $A^{\eta}$ corresponding to $\eta$.
@@
Given these results, a rich theory follows. The set of eigenvalues of the operator $A^{\theta}$ is identical to the set of possible values of $\theta$. The variable $\theta$ is maximal if and only if all eigenvalues are simple. In general, the eigenspaces of $A^{\theta}$ are in one-to-one correspondence with questions ‘What is $\theta$’/ ‘What will $\theta$ be if we measure it?’ together with sharp answers $\theta=u$.
If $\theta$ is maximal as an accessible variable, the eigenvectors of $A^{\theta}$ have a similar interpretation. In my opinion, such eigenvectors, where $\theta$ is some meaningful variable, should be taken as the only possible state vectors. These have straightforward interpretations, and from this version of the theory also a number of so-called ‘quantum paradoxes’ can be illuminated, see Helland (2022b,d).
This not only points at a new foundation of quantum theory, and it also suggests a general epistemic interpretation of the theory: Quantum theory is not directly a theory about the world, but a theory about an actor’s knowledge of the world. Versions of such an interpretation already exist, and it is one of the very many suggested interpretations of quantum mechanics.
@@
To any ket vector $|a;k\rangle$ there corresponds a bra vector $\langle a;k|$, the corresponding complex conjugate row vector. Scalar products between two state vectors can then be written $\langle a;k|b;j\rangle$. I will also need scalar products of the form $\langle a;k|A|a;k\rangle$, where $A$ is an operator.
@@ 2.2 A simple model of a person’s decisions
Consider a person $C$ in some decision situation. Say that he has the choice between a finite set of actions $a_{1},...,a_{r}$. Relative to this situation we can define a finite-valued decision variable $\theta$, taking the different values $1,...,r$, such that $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$. If $C$ really is able to make a decision here and carry out the actions, we say that $\theta$ is an accessible variable. In analogy with the situation in the previous subsection, the variable $\theta$ is in correspondence with a person $C$, in fact, here $\theta$ belongs to the mind of $C$.
@@
This must be made precise. A decision problem is said to be maximal if $C$ is just able to make up his mind with respect to this decision; if the problem is made slightly more complicated, he is not able to make a decision. Let two different maximal, decision variables be $\theta$ and $\eta$, where $\theta=i$ corresponds to the action $a_{i}$ $(i=1,...,r)$, and $\eta=j$ corresponds to the action $b_{j}$ $(j=1,...,r)$. Then, by the theory of the previous subsection, we can model the situation by using quantum theory.
Note that this simple model does not cover all situations. Sometimes we have a choice between an infinite number of possibilities, and sometimes the outer context changes during the decision process. Nevertheless, the simple model is a good starting point.
@@ 2.3 The Hilbert space formalism for decisions
It is well known that our minds may be limited, for instance when faced with difficult decisions. We will first mention a side result in this direction from the present development.
In Helland (2022c), Theorem 2 says essentially: Imagine a person $C$ which in some context has two related maximal accessible variables $\theta$ and $\eta$ in his mind. Impose a specific symmetry assumption. Then $C$ cannot simultaneously have in mind any other maximal accessible variable which is related to $\theta$, but not related to $\eta$. It was claimed by Helland (2022c) that the violation of a famous inequality by practical Bell experiments can be understood on the basis of this theorem. See also Helland (2023c).
Note that this result has the qualification ‘at the same time’, and indicates a specific restriction to two maximal variables. But the human mind is very flexible. Taking time into account, we can think of very many variables, even ones that are not related.
For the present article, however, the direct results from Helland (2022b, 2023b) are more important. Consider again a decision situation, and assume the simple model of the previous Subsection. In particular let $C$ at the same time be confronted with at least two different maximal related decision processes. Then the following hold:
- Each decision variable $\theta$ is associated with a self-adjoint operator $A^{\theta}$, whose eigenvalues are that possible values of $\theta$.
@@
- In the maximal case, the eigenvectors of the operator can be given interpretation: They are coupled to one particular decision process and a specific choice in this decision process: In concrete terms, the eigenvectors $v$ are in one-to-one correspondence with 1) some maximal accessible decision variable $\theta$, and 2) a specific value $u$ of $\theta$. In other words, the possible eigenvectors are in one-to-one correspondence with 1) the question ‘Which decision process?’ and 2) ‘Which action did this decision process lead to?’.
@@
This can be taken as a starting point of quantum decision theory, but to develop this theory further, we need to be able to calculate probabilities of the various decisions.
@@ 2.4 The Born rule; general arguments
where the pure state $|b;j\rangle$ corresponds to the event $\theta^{b}=u^{b}_{j}$, and where $\mathbf{u}^{b}=(u^{b}_{1},...,u^{b}_{r})$ is a vector of actual values of the parameter $\theta^{b}$.
The interpretation of the likelihood effect $F^{b}(z^{b},\tau)$ can be formulated as follows: (1) We have posed an inference question on the accessible conceptual variable $\theta^{b}$, a question that may consist of several subquestions. (2) We have specified the relevant likelihood for the data. The question itself and the likelihood for all possible answers to the question, formulated in terms of state vectors, can be recovered from the likelihood effect.
The likelihood effect is closely connected to the concept of an operator-valued measure; see a discussion in Helland (2021). Since the focused question assumes discrete data, each likelihood is in the range $0\leq p\leq 1$. In the quantum mechanical literature, an effect is any operator with eigenvalues in the range $[0,1]$.
@@
Assumption 1 Consider in the context $\tau$ an epistemic setting where the likelihood principle from statistics is satisfied, and the whole situation is observed by an actor $C$ whose decisions can be modelled to be influenced by a superior being $D$. Assume that $D$’s probabilities for the situation are $q$, and that $D$ can be seen to be perfectly rational in agreement with the Dutch Book Principle.
@@
A situation where Assumption 1 holds, will be called a rational epistemic setting. It will be seen to imply essential aspects of quantum mechanics. Below we will discuss whether or not it also can be coupled to certain macroscopic situations, in particular decision situations.
In Helland (2021), a generalized likelihood principle is proved from the ordinary likelihood principle: Given some experiment, or more generally, some context $\tau$ connected to an experiment, any experimental evidence will under weak assumptions be a function of the likelihood effect $F$. In particular, the probability $q$ is a function of $F$: $q(F|\tau)$.
Theorem 1 Assume a rational epistemic setting, and assume a fixed context $\tau$. Let $F_{1}$ and $F_{2}$ be two likelihood effects in this setting, and assume that $F_{1}+F_{2}$ also is an effect. Then the experimental evidence, taken as the epistemic probabilities related to the data of the performed experiments, satisfy
@@
The further derivations rely on a very elegant theorem by Busch (2003): Let in general $\mathcal{H}$ be any separable Hilbert space. Recall that an effect $F$ is any operator on the Hilbert space with eigenvalues in the range $[0,1]$. A generalized probability measure $\mu$ is a function of the effects with the properties
@@
It is now easy to see that $q(F|\tau)$ on the likelihood effects of the previous Section is a generalized probability measure if Assumption 1 holds: (1) follows since $q$ is a probability; (2) since $F=I$ implies that the likelihood is 1 for all values of the e-variable; finally (3) is a consequence of the corollary of Theorem 1. Hence there is a density operator $\rho=\rho(\tau)$ such that $p(z|\tau)=\mathrm{trace}(\rho(\tau)F)$ for all ideal likelihood effects $F=F(z)$. This is a result that is valid for all experiments, and can be seen as a first general version of Born’s formula.
@@
Define now a perfect experiment as one where the measurement uncertainty can be disregarded. The quantum mechanical literature operates very much with perfect experiments that result in well-defined states $|j\rangle$. From the point of view of statistics, if, say the 99% confidence or credibility region of $\theta^{b}$ is the single point $u_{j}^{b}$, we can infer approximately that a perfect experiment has given the result $\theta^{b}=u_{j}^{b}$.
@@
An advantage of using the version of Gleason’s theorem due to Busch (Theorem 2) in the derivation of Born’s formula, is that this version is valid also in dimension 2. Other derivations using the same point of departure, are Caves et al. (2004) and Wright and Weigert (2019). In Wright and Weigert (2021) the class of general probabilistic theories which also admits Gleason-type theorems is identified. But for instance, Auffeves and Granger (2019) derive the Born formula from other postulates.